OpenStudy (anonymous):

Find the sum of the infinite geometric series 5+ 5/3+ 5/9+ 5/27...

OpenStudy (anonymous):

15/2

jimthompson5910 (jim_thompson5910):

first term is a = 5

jimthompson5910 (jim_thompson5910):

divide any term by the previous term to find the common ratio r r = (5/3) divided by 5 r = (5/3) times (1/5) r = (5*1)/(3*5) r = 5/15 r = 1/3 since |r| < 1 is true, this means we can use the formula \[\Large S = \frac{a}{1-r}\]

OpenStudy (anonymous):

@jim_thompson5910 is the answer 7.5?